# Online Multi-modal Root Cause Analysis

### v1

@@ Abstract
Root Cause Analysis (RCA) is essential for pinpointing the root causes of failures in microservice systems. Traditional data-driven RCA methods are typically limited to offline applications due to high computational demands, and existing online RCA methods handle only single-modal data, overlooking complex interactions in multi-modal systems. In this paper, we introduce OCEAN, a novel online multi-modal causal structure learning method for root cause localization. OCEAN employs a dilated convolutional neural network to capture long-term temporal dependencies and graph neural networks to learn causal relationships among system entities and key performance indicators. We further design a multi-factor attention mechanism to analyze and reassess the relationships among different metrics and log indicators/attributes for enhanced online causal graph learning. Additionally, a contrastive mutual information maximization-based graph fusion module is developed to effectively model the relationships across various modalities. Extensive experiments on three real-world datasets demonstrate the effectiveness and efficiency of our proposed method.
1 Introduction
@@
Root Cause Analysis (RCA) is crucial for identifying the underlying causes of system failures and ensuring the high performance of microservice systems (Wang et al., 2023a; Li et al., 2021; Wang et al., 2023c). Traditional manual root cause analysis is labor-intensive, costly, and error-prone, given the complexity of microservice systems and the extensive volume of data involved. Consequently, effective and efficient root cause analysis methods are vital for pinpointing failures in complex microservice systems and mitigating potential financial losses when system faults occur.
Previous studies in data-driven RCA, particularly those utilizing causal discovery techniques, have primarily focused on constructing causal or dependency graphs (Ikram et al., 2022; Lu et al., 2017; Li et al., 2021; Soldani & Brogi, 2022; Wang et al., 2023c; Zheng et al., 2024a). These graphs depict the causal links between different system entities and key performance indicators (KPIs), thereby enabling the tracing of underlying causes through these structures. For instance, Wang et al. (Wang et al., 2023c) developed a hierarchical graph neural network method that automatically identifies causal relationships both within and between networks to help pinpoint root causes.
Despite significant advances, most of these approaches are designed for offline use and face challenges with real-time implementation in microservice systems due to high computational demands. To address this, Wang et al. (Wang et al., 2023a) introduced an online RCA method that decouples state-invariant and state-dependent information and incrementally updates the causal graph. Li et al. (Li et al., 2022) developed a causal Bayesian network that leverages system architecture knowledge to mitigate potential biases toward new data. However, these online RCA methods are limited to handling single-modal data.
@@
Recently, multi-modal data, such as system metrics and logs, are commonly collected from microservice systems, revealing the complex nature of system failures (Zheng et al., 2024a). For instance, failures such as “Database Query Failures” might be overlooked if only system metrics are considered, whereas issues like “Disk Space Full” are more effectively identified through combined analysis of metrics and logs. This underscores the importance of using multi-modal data for a thorough understanding of system failures. By integrating information from various sources, we can detect the abnormal patterns of system failures that might not be evident when analyzing single-modal data.
@@
To bridge this gap, this paper aims to propose an online multi-modal causal structure learning method for identifying root causes in microservice systems. Formally, given the system KPI data along with multi-modal microservice data including metrics and log data, our objective is to develop an online multi-modal causal graph that identifies the top $k$ system entities most relevant to the system KPI. Three major challenges exist in this task. (C1) Capturing Long-term Temporal Dependencies: Current auto-regressive based RCA methods (Wang et al., 2023a; Zheng et al., 2024a) are limited to capturing short-term temporal dependencies. However, some system faults, such as Distributed Denial of Service (DDoS) attacks, may persist for extended periods. Effectively capturing these long-term temporal dependencies is crucial for identifying various types of system faults. (C2) Capturing the Correlation of Multi-dimensional Factors: Existing RCA approaches (Ikram et al., 2022; Wang et al., 2023c; Zheng et al., 2024a) often analyze abnormal patterns from multiple factors individually, such as CPU usage or memory usage from system metrics and frequency or golden signal from system logs, overlooking potential relationships among these factors from both modalities. Furthermore, these methods often consider all factors as equally important; however, in real applications, certain factors prove to be considerably more crucial than others. It is vital, therefore, to reassess the contributions of each factor to the learning of causal structures. (C3) Learning Multi-modal Causal Structures: Effectively capturing the relationships between different modalities in an online setting is crucial. Simply combining causal graphs from individual modalities can be problematic, especially if one modality is of lower quality.
@@
To tackle these challenges, we introduce OCEAN, Online Multi-modal Causal Structure LEArNing, for root cause identification in microservice systems. Specifically, we propose to encode long-term temporal dependencies using a dilated convolutional neural network (Yu & Koltun, 2016) and forecast future values based on the $p$-th lagged historical data. We further develop a multi-factor attention mechanism to analyze the correlations among various factors and reassess their importance for causal graph learning. Additionally, we propose a contrastive mutual information estimation technique to model the relationships of different modalities. Our contributions can be summarized as follows:
@@
  • We propose employing a dilated convolutional neural network to capture long-term temporal dependencies and graph neural networks to model causal relations among system entities.
  • We design a multi-factor attention mechanism to analyze the relationships among different factors and reassess their impact on online causal graph learning.
@@
2 Preliminary and Related Work
@@
Key Performance Indicator (KPI). In a microservice system, KPIs serve as invaluable metrics for assessing the effectiveness and productivity of the architecture (Podgórski, 2015). They play an indispensable role in monitoring and managing different aspects of microservices to uphold optimal performance levels. Common KPIs encompass latency and service response time. High values in these metrics typically indicate suboptimal system performance or potential system failure.
@@
Entity Metrics. Entity metrics are the measurable time-series attributes that provide insights into the performance and status of services within a system (Bogner et al., 2017). These entities encompass various components such as physical machines, containers, virtual machines, and pods. In microservice architectures, typical entity metrics include CPU utilization, memory usage, disk I/O activity, packet transmission rate, and etc. These metrics are extensively employed to detect anomalous behavior and pinpoint potential causes of system failures in microservice environments (Wang et al., 2023a; b; Zheng et al., 2024a; Soldani & Brogi, 2023; Liu et al., 2021).
Root Cause Analysis. Current root cause analysis (RCA) methods can be categorized into two main branches: single-modal RCA methods and multi-modal RCA methods. Single-modal RCA methods primarily investigate causal relationships among system components using one type of data only (Sporleder et al., 2019; Duan et al., 2020; Meng et al., 2020; Soldani & Brogi, 2022; Aggarwal et al., 2020; Li et al., 2021). For instance, Liu et al. (Liu et al., 2021) generate a service call graph based on domain-specific software and rules, while Wang et al. (Wang et al., 2023c) construct causal networks from time series data. However, these methods often exhibit suboptimal performance due to their reliance on single-modal data. To enhance accuracy, recent research integrates multi-modal data for RCA (Yu et al., 2023; Hou et al., 2021; Zheng et al., 2024a). Nezha (Yu et al., 2023) and PDiagnose (Hou et al., 2021) extract and combine information from each modality individually, while MULAN (Zheng et al., 2024a) and MM-DAG (Lan et al., 2023) consider interactions among modalities, constructing comprehensive causal graphs. Despite notable progress, these approaches are implemented offline, necessitating extensive data collection and retraining for new faults. Wang et al. (Wang et al., 2023a) enable online root cause identification by decoupling state-invariant and state-dependent information to learn a causal graph for root cause identification. However, their focus remains on single-modal data. Recently, large language model (LLM)-based approaches have emerged as a new research direction for learning causal relations in root cause identification, owing to the success of LLMs in tackling complex tasks (Chen et al., 2024; Shan et al., 2024; Goel et al., 2024; Zhou et al., 2024; Roy et al., 2024; Wang et al., 2024). For example, Chen et al. (Chen et al., 2024) introduce RCACopilot, an on-call system powered by LLMs to automate RCA for cloud incidents. Similarly, Shan et al. (Shan et al., 2024) propose an approach that first identifies log messages indicating configuration-related errors, then localizes suspected root-cause configuration properties based on these log messages and LLM-generated configuration settings. While LLMs could effectively learn temporal dependencies within the metric data of individual system entities, they often struggle to capture interdependencies—such as causal relationships between different system entities—leading to higher computational costs. Unlike existing RCA methods, this paper addresses the online multi-modal RCA problem by uniquely modeling long-term temporal dependencies while simultaneously capturing the cross-modal correlation of multiple factors.
Table 1: Notation Table
$\bm{X}_{M}^{0}$ the historical metric data
$\bm{X}_{M}^{i}$ the $i$-th batch of the system metric
$\bm{X}_{L}^{0}$ the historical log data
$\bm{X}_{L}^{i}$ the $i$-th batch of system log
$T_{1}$ the length of the historical metric time-series data
@@
$n-1$ the number of system entities
@@
3 Methodology
In this section, we first present the problem statement and then introduce OCEAN, an online causal structural learning method designed to identify root causes using multi-modal data. We propose three modules to tackle the challenges outlined in the introduction: long-term temporal causal structure learning, multi-factor attention mechanism, and contrastive multi-modal learning. Subsequently, we identify potential root causes through the network propagation-based root cause identification module and establish stopping criteria. The overview of the proposed OCEAN is provided in Figure 1.
Figure 1: The overview of the proposed framework OCEAN with four main modules: long-term temporal causal structure learning, representation learning with multi-factor attention, graph fusion with contrastive multi-modal learning, and network propagation-based root cause localization.
3.1 Problem Statement
Let $\mathcal{X}_{M}=\{\bm{X}^{0}_{M},\bm{X}^{1}_{M},...,\bm{X}^{\mathcal{T}}_{M}\}$ represent $\mathcal{T}+1$ multi-variate time series data for entity metrics. Here, $\bm{X}_{M}^{0}$ is the historical metric data, and $\bm{X}_{M}^{i}$, $i\in[1,\ldots,\mathcal{T}]$, is the $i^{\textrm{th}}$ batch for the metric data, with $T_{1}$ denoting the length of historical metric data, $T_{2}$ the length of each batch, $n-1$ the number of system entities, and $d_{M}$ the number of different metrics. Similarly, $\mathcal{X}_{L}=\{\bm{X}^{0}_{L},\bm{X}^{1}_{L},...,\bm{X}^{\mathcal{T}}_{L}\}$ represents $\mathcal{T}+1$ multi-variate time series data for system logs. Assuming preprocessing has converted the logs into multi-variate time series data (details of converting log data into time series can be found in Appendix C), $\bm{X}_{L}^{0}$ is the historical log data, and $\bm{X}_{L}^{i}$, $i\in[1,\ldots,\mathcal{T}]$, is the $i^{\textrm{th}}$ batch for system logs, where $d_{L}$ is the number of different log attributes/features. The system KPI is denoted as $\mathbf{y}=\{\bm{y}^{0},\bm{y}^{1},...,\bm{y}^{\mathcal{T}}\}$, with $\bm{y}^{0}$ and $\bm{y}^{i}$, $i\in[1,\ldots,\mathcal{T}]$, representing KPI data with lengths $T_{1}$ and $T_{2}$, respectively. Our goal is to construct a causal graph $\mathcal{G}=\{\mathcal{V},\mathcal{A}\}$ to identify the top $k$ system entities most relevant to $\mathbf{y}$. Here, $\mathcal{V}$ represents the set of vertices, $\mathcal{A}\in\mathbb{R}^{n\times n}$ denotes the adjacency matrix, and $n$ is the total number of entities plus the system KPI. To simplify, we replicate the KPI $d_{M}$ times to match the number of metrics. This allows us to concatenate the system metric time-series data and KPI, yielding $\hat{\bm{X}}_{M}^{0}\in\mathbb{R}^{n\times d_{M}\times T_{1}}$ and $\hat{\bm{X}}_{M}^{i}\in\mathbb{R}^{n\times d_{M}\times T_{2}}$ for metric data. Similarly, we combine the system log time-series data and KPI, denoted as $\hat{\bm{X}}_{L}^{0}\in\mathbb{R}^{n\times d_{L}\times T_{1}}$ and $\hat{\bm{X}}_{L}^{i}\in\mathbb{R}^{n\times d_{L}\times T_{2}}$. Online RCA is a crucial step in system diagnosis post online fault/failure detection, which falls beyond the scope of this work. Here, we employ the Multivariate Singular Spectrum Analysis (MSSA) model (Alanqary et al., 2021), a state-of-the-art method for online failure detection, to identify the triggers for the root cause analysis process. We summarize the notation in Table 1.
3.2 Long-term Temporal Causal Structure Learning
To capture temporal causal relations among various system entities and KPIs, the Vector Autoregression Model (VAR) (Stock & Watson, 2001) is often employed (Wang et al., 2023a; b; Zheng et al., 2024a) due to its effectiveness in capturing dynamic interactions among variables in time series data. Specifically, given the two-way matrix $\bm{X}_{M,i}\in\mathbb{R}^{n\times T_{1}}$ for the $i^{\textrm{th}}$ system metric, our objective is to minimize a VAR-based loss function as follows:
$\hat{\bm{X}}_{M,i}^{t}$ $=\bm{W}^{1}\bm{X}_{M,i}^{t-1}+\bm{W}^{2}\bm{X}_{M,i}^{t-2}+\cdot\cdot\cdot+\bm{W}^{t-1}\bm{X}_{M,i}^{1}+\epsilon$
$\mathcal{L}_{var}$ $=\sum_{i=1}^{d_{M}}||\bm{X}^{t}_{M,i}-F(\hat{\bm{X}}^{t}_{M,i},\mathcal{A},\theta)||^{2}$ (1)
where $\hat{\bm{X}}_{M,i}^{t}$ represents the prediction for the $i^{\textrm{th}}$ system metric, $\bm{W}^{i}\in\mathbb{R}^{n\times n}$ denotes the weight matrix, $\epsilon\in\mathbb{R}^{n}$ signifies the error variable, $F(\cdot)$ represents the graph neural network (Kipf & Welling, 2017) parameterized by $\theta$, $\mathcal{A}$ is the learnable causal graph capturing the relationships among node entities and KPIs, and $\bm{X}_{M,i}^{t}$ denotes the future value.
This model is also known as the $t^{\textrm{th}}$ order VAR-based model, where $t$ defines the range of temporal dependencies it can capture. However, a recent study (Lin et al., 2020) indicates that as the time lag $t$ increases, the autoregressive model becomes computationally expensive, making it challenging to capture long-term temporal dependencies in online settings. Similarly, many existing temporal modeling techniques, such as Recurrent Neural Networks (RNNs) and Transformers (Vaswani et al., 2017), also incur high computational costs, limiting their applicability for online root cause analysis (see Appendix A for detailed discussion).
To address this issue, we propose a module based on dilated convolution and graph neural networks to efficiently capture the long-term temporal dependencies and causal relations among system entities and KPIs. Different from the VAR-based methods (Wang et al., 2023a; b; Zheng et al., 2024a) that take as input the 2-way matrix, we propose to capture the long-term temporal dependency via 3-way tensors. Specifically, given four 3-way tensors (i.e., historical metric data $\hat{\bm{X}}_{M}^{0}$, historical log data $\hat{\bm{X}}_{L}^{0}$, the current batch of metric data $\hat{\bm{X}}_{M}^{i}$ and log data $\hat{\bm{X}}_{L}^{i}$), we follow the idea of LSTM (Hochreiter & Schmidhuber, 1997) and Gated Temporal Convolutional Network (TCN) (Wu et al., 2019) to model the temporal dependency for the historical and current batches of time series for two modalities as follows:
$g(\bm{x},\bm{f})$ $=\bm{x}*\bm{f}=\sum_{\tau=0}^{K-1}\bm{f}(\tau)\cdot\bm{x}(t-d\times\tau)$ (2)
$\bm{H}_{v}^{0}$ $=\tanh(g(\hat{\bm{X}}_{v}^{0},\bm{f}_{1}))\odot\sigma_{1}(g(\hat{\bm{X}}_{v}^{0},\bm{f}_{2}))$ (3)
$\bm{H}_{v}^{i}$ $=\tanh(g(\hat{\bm{X}}_{v}^{i},\bm{f}_{3}))\odot\sigma_{1}(g(\hat{\bm{X}}_{v}^{i},\bm{f}_{4}))$ (4)
$\hat{\bm{O}}_{v}^{0}$ $=MFL(\hat{\bm{H}}_{v}^{0}),\hat{\bm{O}}_{v}^{i}=MFL(\hat{\bm{H}}_{v}^{i})$ (5)

### v2

# Online Multi-modal Root Cause Identification in Microservice Systems

@@ Abstract
Root Cause Analysis (RCA) is essential for pinpointing the root causes of failures in microservice systems. Traditional data-driven RCA methods are typically limited to offline applications due to high computational demands, and existing online RCA methods handle only single-modal data, overlooking complex interactions in multi-modal systems. In this paper, we introduce OCEAN, a novel online multi-modal causal structure learning method for root cause localization. OCEAN introduces a long-term temporal causal learning module with two encoders: one captures stable causal dependencies from historical data, while the other models short-term variations in the current batch data. We further design a multi-factor attention mechanism to analyze and reassess the relationships among different metrics and log indicators/attributes for enhanced online causal graph learning. Additionally, a contrastive mutual information maximization-based graph fusion module is developed to effectively model the relationships across various modalities. Extensive experiments on three real-world datasets demonstrate the effectiveness and efficiency of our proposed method.
I Introduction
@@
Root Cause Analysis (RCA) is crucial for identifying the underlying causes of system failures and ensuring the high performance of microservice systems [wang2023incremental, li2021practical]. Traditional manual root cause analysis is labor-intensive, costly, and error-prone, given the complexity of microservice systems and the extensive volume of data involved. Consequently, effective and efficient root cause analysis methods are vital for pinpointing failures in complex microservice systems and mitigating potential financial losses when system faults occur. Prior works leveraging causal discovery have focused on constructing causal or dependency graphs [DBLP:conf/nips/IkramCMSBK22, lu2017log, li2021practical], which reveal causal links among system entities and key performance indicators to trace underlying faults.
@@
Despite significant advances, most of these approaches are designed for offline use and face challenges with real-time implementation in microservice systems due to high computational demands. To address this, Wang et al. [wang2023incremental] introduced an online RCA method that decouples state-invariant and state-dependent information and incrementally updates the causal graph. Li et al. [DBLP:conf/kdd/0005LYNZSP22] developed a causal Bayesian network that leverages system architecture knowledge to mitigate potential biases toward new data. However, these online RCA methods are limited to handling single-modal data. Recently, multi-modal data, such as system metrics and logs, are commonly collected from microservice systems, revealing the complex nature of system failures [DBLP:conf/www/ZhengCHC24]. For instance, failures such as “Database Query Failures” might be overlooked if only system metrics are considered, whereas issues like “Disk Space Full” are more effectively identified through combined analysis of metrics and logs. This underscores the importance of using multi-modal data for a thorough understanding of system failures. By integrating information from various sources, we can detect the abnormal patterns of system failures that is evident when analyzing single-modal data.
@@
To bridge this gap, this paper proposes an online multi-modal causal structure learning method for RCA in microservice systems. Given the system KPI and multi-modal data (metrics and log data), our objective is to construct an online multi-modal causal graph that identifies the top $k$ system entities most relevant to the system KPI. Three major challenges arise in this task: (C1) Enabling Long-term Temporal Causal Relationship Learning in the Online Setting: Existing temporal/sequential modeling techniques [chen2024automatic, shan2024face, zhou2024causalkgpt, wang2023incremental, DBLP:conf/www/ZhengCHC24], (e.g., RNNs, Transformers, and LLMs) are computationally expensive and limited to short-term dependencies. However, some system faults, such as Distributed Denial of Service (DDoS) attacks, may persist for extended periods. Effectively capturing these long-term temporal dependencies in the online setting is crucial for efficiently identifying various types of system faults. (C2) Capturing the Correlation of Multi-dimensional Factors: Existing RCA approaches [DBLP:conf/nips/IkramCMSBK22, wang2023interdependent, DBLP:conf/www/ZhengCHC24] often analyze abnormal patterns from multiple factors individually, such as CPU usage or memory usage from system metrics and frequency or golden signal from system logs, overlooking potential relationships among these factors from both modalities. Furthermore, these methods often consider all factors as equally important; however, in real applications, certain factors prove to be considerably more crucial than others. It is vital, therefore, to reassess the contributions of each factor to the learning of causal structures. (C3) Learning Multi-modal Causal Structures: Effectively capturing the relationships between different modalities [DBLP:conf/www/ZhouZZLH20, DBLP:conf/acl/0006ZJFJBH025, DBLP:journals/corr/abs-2502-08942, DBLP:journals/corr/abs-2504-07394, DBLP:conf/kdd/ZhengJLTH24, DBLP:journals/corr/abs-2102-07751, DBLP:conf/sdm/ZhengCH19] in an online setting is crucial. Simply combining causal graphs from individual modalities can be problematic, especially if one modality is of lower quality.
@@
To tackle these challenges, we introduce OCEAN, Online Multi-modal Causal Structure LEArNing, for root cause identification in microservice systems. Specifically, we design a long-term temporal causal relationship learning module with two encoders, one for capturing the long-term temporal dependencies and invariant causal relation for historical data and another for modeling the small changes within current batch data. We further develop a multi-factor attention mechanism to analyze the correlations among various factors and reassess their importance for causal graph learning. Additionally, we propose a contrastive mutual information estimation technique to model the relationships of different modalities. Our contributions can be summarized as follows:
@@
  • We propose long-term temporal causal relationship learning module with two encoders, aiming to efficiently capture long-term temporal dependencies and causal relations for both historical data and current batch data.
@@
II Preliminary and Related Work
@@
Key Performance Indicator (KPI). In a microservice system, KPIs serve as invaluable metrics for assessing the effectiveness and productivity of the architecture [podgorski2015measuring]. They play an indispensable role in monitoring and managing different aspects of microservices to uphold optimal performance levels. Common KPIs encompass latency and service response time. High values in these metrics typically indicate suboptimal performance or potential failure.
@@
Entity Metrics. Entity metrics are the measurable time-series attributes that provide insights into the performance and status of services within a system [bogner2017automatically]. These entities encompass various components such as physical machines, containers, virtual machines, and pods. In microservice architectures, typical entity metrics include CPU utilization, memory usage, disk I/O activity, packet transmission rate, and etc. These metrics are extensively employed to detect anomalous behavior and pinpoint potential causes of system failures in microservice environments [DBLP:conf/www/ZhengCHC24, soldani2022anomaly].
Root Cause Analysis. Current root cause analysis (RCA) methods can be categorized into two main branches: single-modal RCA methods and multi-modal RCA methods. Single-modal RCA methods primarily investigate causal relationships among system components using one type of data only [soldani2022anomaly, li2021practical, wang2023interdependent, wang2023incremental], while multi-modal RCA methods [yu2023nezha, DBLP:conf/ispa/HouJWLH21, DBLP:conf/www/ZhengCHC24] benefit from leveraging the rich data sources to achieve better performance. Recently, large language model (LLM)-based approaches have emerged as a new research direction for learning causal relations in root cause identification, owing to the success of LLMs in tackling complex tasks [chen2024automatic, shan2024face, goel2024x, zhou2024causalkgpt, roy2024exploring, wang2023rcagent]. Unlike existing RCA methods, this paper addresses the online multi-modal RCA problem by uniquely modeling long-term temporal dependencies while simultaneously capturing the cross-modal correlation of multiple factors.
III Methodology
In this section, we first present the problem statement and then introduce OCEAN, an online causal structural learning method designed to identify root causes using multi-modal data. We propose three modules to tackle the challenges outlined in the introduction: long-term temporal causal relationship learning, contrastive multi-modal learning and network propagation-based root cause identification module. The overview of OCEAN is provided in Figure 1.
Figure 1: The overview of the proposed framework OCEAN with three main modules: long-term temporal causal relationship learning, graph fusion with contrastive multi-modal learning, and network propagation-based root cause localization. Both static encoder and dynamic encoder are trained in the offline stage but only dynamic encoder will be finetuned with batch (streaming) data during the online stage.
III-A Problem Statement
Let $\mathcal{X}_{M}=\{\mathbf{X}_{M}^{0},\mathbf{X}_{M}^{1},\ldots,\mathbf{X}_{M}^{\mathcal{T}}\}$ denote the multivariate time-series data of system metrics across n1n-1 entities, where $\mathbf{X}_{M}^{0}\in\mathbb{R}^{(n-1)\times d_{M}\times T_{1}}$ represents a large historical dataset and $\mathbf{X}_{M}^{i}\in\mathbb{R}^{(n-1)\times d_{M}\times T_{2}}$ for $i\in[1,\mathcal{T}]$ are smaller sequential streaming batches. Here, $d_{M}$ is the number of metric features, and $T_{1}$ and $T_{2}$ denote the temporal lengths of the historical and streaming segments, respectively. Similarly, $\mathcal{X}_{L}=\{\mathbf{X}_{L}^{0},\mathbf{X}_{L}^{1},\ldots,\mathbf{X}_{L}^{\mathcal{T}}\}$ denotes the multivariate time-series data of system logs, where each $\mathbf{X}_{L}^{i}\in\mathbb{R}^{(n-1)\times d_{L}\times T}$ is obtained by preprocessing log records into structured time-series format, and $d_{L}$ is the number of log features. The system KPI is represented as $\mathbf{y}=\{\mathbf{y}^{0},\mathbf{y}^{1},\ldots,\mathbf{y}^{\mathcal{T}}\},$ where $\mathbf{y}^{0}\in\mathbb{R}^{T_{1}}$ and $\mathbf{y}^{i}\in\mathbb{R}^{T_{2}}$ correspond to historical and batch KPI data, respectively. We consider $n$ nodes in total: n1n-1 microservice entities and one KPI node. The causal structure among them is modeled by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{A}),$ where $\mathcal{V}$ denotes the set of vertices and $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, with $\mathcal{A}_{ij}\neq 0$ indicating a causal influence from node $i$ to node $j$. For unified processing, we replicate the KPI $d_{M}$ times and concatenate it with the metric and log data, forming $\hat{\mathbf{X}}_{M}^{0}\in\mathbb{R}^{n\times d_{M}\times T_{1}},\quad\hat{\mathbf{X}}_{M}^{i}\in\mathbb{R}^{n\times d_{M}\times T_{2}},$ and $\hat{\mathbf{X}}_{L}^{0}\in\mathbb{R}^{n\times d_{L}\times T_{1}},\quad\hat{\mathbf{X}}_{L}^{i}\in\mathbb{R}^{n\times d_{L}\times T_{2}}.$ In the offline setting, only the historical data ($\mathbf{X}^{0}$, $\mathbf{y}^{0}$) are available for model training, while in the online setting, streaming batches ($\mathbf{X}^{i}$, $\mathbf{y}^{i}$) are also used for root cause analysis. Finally, we employ the Multivariate Singular Spectrum Analysis (MSSA) model [alanqary2021change], a state-of-the-art method for online anomaly detection, to trigger the root cause analysis process. We summarize the notations in Table I.
TABLE I: Notation Table
$\bm{X}_{M}^{0}$ the historical system metric (offline data)
$\bm{X}_{M}^{i}$ the $i$-th batch of the system metric (streaming data)
$\bm{X}_{L}^{0}$ the historical system log (offline data)
$\bm{X}_{L}^{i}$ the $i$-th batch of system log (streaming data)
$T_{1}$ the length of the historical metric data
@@ III-A Problem Statement
n1n-1 the number of system entities
@@
III-B Long-term Temporal Causal Relationship Learning
To effectively capture temporal causal relationships among system entities and KPIs, we adopt the Vector Autoregression (VAR) model [stock2001vector], a classical yet powerful framework for modeling dynamic interactions in multivariate time series. Given the two-dimensional matrix $\mathbf{X}_{M,i}\in\mathbb{R}^{n\times T_{1}}$ representing the $i^{\textrm{th}}$ system metric, the model predicts future values as:
$\hat{\mathbf{X}}_{M,i}^{t}$ $=\mathbf{W}^{1}\mathbf{X}_{M,i}^{t-1}+\mathbf{W}^{2}\mathbf{X}_{M,i}^{t-2}+\cdots+\mathbf{W}^{t-1}\mathbf{X}_{M,i}^{1}+\bm{\epsilon},$
$\mathcal{L}_{\text{var}}$ $=\sum_{i=1}^{d_{M}}\|\mathbf{X}_{M,i}^{t}-F(\hat{\mathbf{X}}_{M,i}^{t},\mathcal{A},\theta)\|^{2},$ (1)
where $\hat{\mathbf{X}}_{M,i}^{t}$ denotes the predicted values for the $i^{\textrm{th}}$ metric, $\mathbf{W}^{k}\in\mathbb{R}^{n\times n}$ are lag-specific weight matrices, $\bm{\epsilon}$ is the residual error, and $F(\cdot)$ represents a graph neural network (GNN) [DBLP:conf/iclr/KipfW17] parameterized by $\theta$. The learnable adjacency matrix $\mathcal{A}$ encodes causal relationships among entities and the KPI. Although a $t^{\textrm{th}}$-order VAR model captures long-range dependencies, it becomes computationally expensive as $t$ increases [lin2020limitations]. Similarly, recurrent architectures (e.g., RNNs) and Transformers [DBLP:conf/nips/VaswaniSPUJGKP17] often suffer from high computational overhead, limiting their suitability for online root cause analysis. To address these challenges, we propose a dual-encoder framework consisting of a Static Encoder and a Dynamic Encoder to efficiently learn long-term temporal dependencies and causal structures for historical and streaming data, respectively. The Static Encoder extracts stable, invariant causal relationships among entities in the offline stage, while the Dynamic Encoder incrementally adapts the learned causal structure during online updates.
Static Encoder
In microservice systems, each entity is associated with multiple metric and log indicators (e.g., CPU usage, memory usage, log frequency, golden signals, etc.). Conventional RCA methods typically analyze each factor independently, overlooking cross-factor dependencies and their varying importance under different abnormal patterns. To address this limitation, we design the Static Encoder to jointly model temporal dependencies and multi-factor correlations across two modalities (metrics and logs) using attention-based multi-factor learning [DBLP:conf/nips/LuYBP16, DBLP:conf/nips/VaswaniSPUJGKP17]. Given historical metric data $\hat{\mathbf{X}}_{M}^{0}$ and log data $\hat{\mathbf{X}}_{L}^{0}$, we first capture temporal dependencies via a gated temporal convolutional network (TCN) [DBLP:conf/ijcai/WuPLJZ19]:
$g(\mathbf{x},\mathbf{f})$ $=\mathbf{x}*\mathbf{f}=\sum_{\tau=0}^{K-1}\mathbf{f}(\tau)\cdot\mathbf{x}(t-d\times\tau),$ (2)
$\mathbf{H}_{v}^{0}$ $=\tanh(g(\hat{\mathbf{X}}_{v}^{0},\mathbf{f}_{1}))\odot\sigma(g(\hat{\mathbf{X}}_{v}^{0},\mathbf{f}_{2})),$ (3)